MONT 104Q – Mathematical Journeys Reading/Discussion Questions on Cheryl Strayed, “Wild” December, 2015

1. “The Pacific Crest Trail wasn’t a world to me then. It was an idea, vague and outlandish, full of promise and mystery. Something bloomed inside me as I traced its jagged line with my finger on a map” (p. 4). Why did the PCT capture Strayed’s imagination at that point in her life? What were the motivations for her journey? As she saw them at the start? At the end? Was she prepared for what she would find? How did the journey change her?

2. Strayed also writes that the point of the PCT “had only to do with how it felt to be in the wild. With what it was like to walk for miles for no reason other than to witness the accumulation of trees and meadows, mountains and deserts, streams and rocks, rivers and grasses, sunrises and sunsets” (p. 207). What are the meanings of the title, Wild? How does the sensation expressed in this quotation help Strayed to find her way back into the world beyond the wilderness?

3. We learn part way into the story that “Cheryl Strayed” is, in part, a pen-name: “Nothing fit until one day when the word strayed came into my mind. Immediately, I looked it up in the dictionary and knew it was mine...: to wander from the proper path, to deviate from the direct course, to be lost, to become wild, to be without a mother or father, to be without a home, to move about aimlessly in search of something, to diverge or digress” (p. 96). Did she choose well? Or does this seem somehow contrived or even “cheesy” in a way?

4. Strayed is apparently very open in her description of her own failings and questionable choices (for instance, about her former drug use, her promiscuity, and the way she wrecked the relationship with her ex-husband). Does some of it seem as though it might be “too good to be true,” or invented for the purpose of creating a good story? Even if all of it is literally true, while she expresses remorse for the things she has done wrong, she doesn’t seem to be ashamed of them. Do you think that indicates strength of character? Or is it a flaw?

6. Strayed says her mother’s death “had obliterated me.... I was trapped by her but utterly alone. She would always be the empty bowl that no one could fill” (p. 267). How did being on the PCT on her mother’s fiftieth birthday help Strayed to heal this wound?

7. “I knew that if I allowed fear to overtake me, my journey was doomed. Fear, to a great extent, is born of a story we tell ourselves, and so I chose to tell myself a different story from the one women are told” (p. 51). What does Strayed mean by “the story that women are told?” And how does that relate to what goes before: “the story we tell ourselves?” What do you think the story Cheryl told herself is? Do you think she was too afraid, or not afraid enough? Was she in real danger? When?

8. The people Strayed meets on the trail are almost entirely men. How does this work in her favor? What role does gender play when removed from the usual structure of society? What was it about Strayed that inspired the generosity of so many strangers on the PCT?

9. On her journey, Strayed carries several totems. What does the black feather mean to her? And the POW bracelet? Why does she find its loss (p. 238) symbolic? What role do books and reading play in this often solitary journey? To lighten her load, Strayed burns each book as she reads it. Why doesn’t she burn the Adrienne Rich poetry collection?

10. Strayed’s hiking pack, also known as “Monster,” is one of those real-life objects that also makes a perfect literary metaphor. What are the symbolic meanings expressed through the hiking pack? Are there other objects she takes with her or acquires along the way that take on deeper meanings? How so?

Note: Most of these questions are adapted from a series of discussion questions for book groups developed by Vintage Books, the publisher of Wild.